4. A ground level ozone mixing ratio of 120 ppb is very unhealthy. How does that level compare to the stratospheric ozone layer, where the concentration of ozone is about 5x1012 molecules cm*?
Answer #1

ppb can be in g/ml or g/g

120 ppb Ozone means we have 120 g of ozone per 1 billion ml of air

volume of air = 10^6 ml = 10^6 cm3      ( since 1ml = 1cm3)

Ozone moles = mass / molar mass of ozone = 120g/ ( 48g/mol) = 2.5 mol

Number of ozone molecules = moles x avagadro number

                = 2.5 mol x 6.022 x 10^23 molecules / 1mol

                   = 15.055 x 10^23

Now the ratio of molecules / voluem = ( 15.055 x 10^23) / ( 10^6 cm3) = 1.505 x 10^18 molecules / cm3

Thus concentration of ozone in ground level is higher compared to startospher ozone concentration which is 5 x 10^12 molecules / cm3